# Using LSDB to enable large-scale catalog distribution, cross-matching, and analytics

### v1

@@ Abstract
The Vera C. Rubin Observatory will generate unprecedented amounts of data, including 60 PB of raw data and 30 trillion observed sources, presenting a significant challenge for large-scale and end-user scientific analysis. As part of the LINCC Frameworks Project we are addressing these challenges with the development of the HATS (Hierarchical Adaptive Tiling Scheme) format and analysis package LSDB (Large Scale Database). HATS partitions data adaptively using a hierarchical tiling system to balance the file sizes, enabling efficient parallel analysis. Recent updates include improved metadata consistency, support for incremental updates, and enhanced compatibility with evolving datasets. LSDB complements HATS by providing a scalable, user-friendly interface for large catalog analysis, integrating spatial queries, crossmatching, and time-series tools while utilizing Dask for parallelization. We have successfully demonstrated the use of these tools with datasets such as ZTF and Pan-STARRS data releases on both cluster and cloud environments. We are deeply involved in several ongoing collaborations to ensure alignment with community needs, with future plans for IVOA standardization and support for upcoming Rubin, Euclid and Roman data. We provide our code and materials at lsdb.io.
@@ 1 Motivation
The LINCC (LSST Interdisciplinary Network for Collaboration And Computing) Frameworks Project is an initiative by the LSST Discovery Alliance to enable such analysis. As part of this effort, we have been developing LSDB (Large Scale Database) and HATS (Hierarchical Adaptive Tiling Scheme) format to deliver and enable end-user analysis on 10TB+ catalog datasets.

### v2

@@ Abstract
The Vera C. Rubin Observatory will generate unprecedented amounts of data, including 60 PB of raw data and 30 trillion observed sources, presenting a significant challenge for large-scale and end-user scientific analysis. As part of the LINCC Frameworks Project we are addressing these challenges with the development of the HATS (Hierarchical Adaptive Tiling Scheme) format and analysis package LSDB. HATS partitions data adaptively using a hierarchical tiling system to balance the file sizes, enabling efficient parallel analysis. Recent updates include improved metadata consistency, support for incremental updates, and enhanced compatibility with evolving datasets. LSDB complements HATS by providing a scalable, user-friendly interface for large catalog analysis, integrating spatial queries, crossmatching, and time-series tools while utilizing Dask for parallelization. We have successfully demonstrated the use of these tools with datasets such as ZTF and Pan-STARRS data releases on both cluster and cloud environments. We are deeply involved in several ongoing collaborations to ensure alignment with community needs, with future plans for IVOA standardization and support for upcoming Rubin, Euclid and Roman data. We provide our code and materials at lsdb.io.
@@ 1 Motivation
The LINCC (LSST Interdisciplinary Network for Collaboration And Computing) Frameworks Project is an initiative by the LSST Discovery Alliance to enable such analysis. As part of this effort, we have been developing LSDB and HATS (Hierarchical Adaptive Tiling Scheme) format to deliver and enable end-user analysis on 10TB+ catalog datasets.